There are 1% probability for a hard drive to crash. Therefore, it has two backups and each
having 2% probability to crash, and all three components are independent of each other. The
stored information is lost only in an unfortunate situation when all three devices crash. What is
the probability that information is saved?

Answer by ikleyn(52777)

.

    It is  P = (1-0.01)*(1-0.02)*(1-0.02) = 0.99*0.98*0.98 = 0.951  (approx.)    ANSWER

We take the  COMPLEMENT  partial probabilities and multiply them.

Solved.

--------------

See the lesson
    - Solving probability problems using complementary probability
    - Solving probability problems using complementary probability REVISITED
in this site.